A portable air compressor rated at 1,600 W draws how many amps from a 12.0 V car battery?
1
Expert's answer
2019-05-27T11:14:54-0400

Electricity power P drawn by the compressor is calculated by


"P=V\cdot I"

where V is the voltage on car battery, I is the current through the compressor. Substituting P=1,600 W and V=12.0 volts, we get


"1,600\,W=12.0\,\,volts\cdot I\,amps"

Now we find the current through the compressor


"I=\frac{1,600\,W}{12.0\,\,volts}\approx 133.3\,amps"

So, the electric current through the compressor is 133.3 amps.